# Text-to-CAD Generation Through Infusing Visual Feedback in Large Language Models

### v2

@@ C.3 GPT Baselines
C.4 Additional Statements on Text2CAD Results
@@
We discovered that the discrepancy stems from the model’s sensitivity to the level of detail in the prompt. Text2CAD performs well only with expert-level prompts, which contain step-by-step sketching guidelines. Our prompts, however, do not include this level of detail 555We are concerned that the impact of detailed prompts containing step-by-step instructions and point coordinates is limited, as they may not be feasible in real-life scenarios.. To ensure consistency, in Table 1, we report Text2CAD’s performance based on their expert prompts when computing the metrics they introduced. Specifically, for each item in the test set, CADFusion and GPT-4o’s predictions were generated using our prompts, while Text2CAD’s predictions were generated using the expert prompt for the same item from their prompt base.
@@
By changing the prompt from the expert-level prompt in their database to an intermediate-level prompt, we observe a similar performance drop. This indicates that our prompting method does not degrade Text2CAD’s performance. Instead, it is an limitation stemmed from Text2CAD itself. Our model, using a simplified prompt, outperforms Text2CAD-expert. Given that the expert-level prompt from Text2CAD is too long and too specific to be feasible in the real designing process, we believe that our quantitative advantage over it is significant.
@@
| CADFusion | 83.71 | 81.99 | 89.97 | 92.79 | 45.67 |
C.5 Additional Quantative Results
@@
C.6 Additional Qualitative Results
@@
C.7 Text to Multiple CAD Figures

### v3

@@ D.3 GPT Baselines
D.4 Additional Statements on Text2CAD Results
@@
We discovered that the discrepancy stems from the model’s sensitivity to the level of detail in the prompt. Text2CAD performs well only with expert-level prompts, which contain step-by-step sketching guidelines. Our prompts, however, do not include this level of detail 666We are concerned that the impact of detailed prompts containing step-by-step instructions and point coordinates is limited, as they may not be feasible in real-life scenarios.. To ensure consistency, in Table 1, we report Text2CAD’s performance based on their expert prompts when computing the metrics they introduced. Specifically, for each item in the test set, CADFusion and GPT-4o’s predictions were generated using our prompts, while Text2CAD’s predictions were generated using the expert prompt for the same item from their prompt base.
@@
By changing the prompt from the expert-level prompt in their database to an intermediate-level prompt, we observe a similar performance drop. This indicates that our prompting method does not degrade Text2CAD’s performance. Instead, it is an limitation stemmed from Text2CAD itself. Our model, using a simplified prompt, outperforms Text2CAD-expert. Given that the expert-level prompt from Text2CAD is too long and too specific to be feasible in the real designing process, we believe that our quantitative advantage over it is non-trivial.
@@
| CADFusion | 83.71 | 81.99 | 89.97 | 92.79 | 19.89 |
D.5 Additional Quantative Results
@@
D.6 Additional Qualitative Results
@@
D.7 Text to Multiple CAD Figures
@@ D.8 Failure Cases
Appendix E Change Log
1.
We updated our chamfer distance (CD) measurement results. The original outcome is computed between distributions and the new one is the accurate, pairwise result. This update does not affect the integrity of our work because the former was an upperbound of the latter measure and actually downgraded our model performance.
2.
We revised writing and updated the related work section according to the reviewers’ recommendations.